Find all the complex numbers z such that

z = (6-8j)4/3

Give your answer in cartesian form
Answer #1

Here Z=(6-8j)4/3

Z=(X+jY) which is complex form

Hence Z=(X+jY) = (8 - j32/3)

So X=8 & Y= - 32/3

Z=X+jY= (8 - j32/3) which is a Cartesian form.